Find the present value of the following future amount. $9000 at 3% compounded semiannually for 7 years What is the present value?
Answer #1

Present value = FV/(1+r/n)nt

Future Valu (FV) = 9,000

r = 3%

n = 2

t = 7

PV = 9,000/(1 + 0.03/2)14

= 9,000/1.23175573

= 7,306.6435

Therefore, Present value is $7,306.6435